### v1

# PhysicsCorrect: A Training-Free Approach for Stable Neural PDE Simulations

Abstract
Neural networks have emerged as powerful surrogates for solving partial differential equations (PDEs), offering significant computational speedups over traditional methods. However, these models suffer from a critical limitation: error accumulation during long-term rollouts, where small inaccuracies compound exponentially, eventually causing complete divergence from physically valid solutions. We present PhysicsCorrect, a training-free correction framework that enforces PDE consistency at each prediction step by formulating correction as a linearized inverse problem based on PDE residuals. Our key innovation is an efficient caching strategy that precomputes the Jacobian and its pseudoinverse during an offline warm-up phase, reducing computational overhead by two orders of magnitude compared to standard correction approaches. Across three representative PDE systems – Navier-Stokes fluid dynamics, wave equations, and the chaotic Kuramoto-Sivashinsky equation – PhysicsCorrect reduces prediction errors by up to 100× while adding negligible inference time (under 5%). The framework integrates seamlessly with diverse architectures including Fourier Neural Operators, UNets, and Vision Transformers, effectively transforming unstable neural surrogates into reliable simulation tools that bridge the gap between deep learning’s computational efficiency and the physical fidelity demanded by practical scientific applications.
1 Introduction
Figure 1: PhysicsCorrect stabilizes neural PDE solver rollouts by projecting erroneous predictions back onto the manifold of physically consistent solutions.
Simulating physical systems governed by partial differential equations (PDEs) is fundamental to numerous scientific and engineering disciplines. Achieving stable long-term rollouts is especially critical for applications such as optimal control, inverse design, and computational imaging. While classical numerical methods like finite-difference (smith1985numerical, ), finite-element (reddy1993introduction, ), and spectral-element methods (patera1984spectral, ) provide accurate solutions, they often demand substantial computational resources, limiting their real-time applicability.
Neural PDE solvers have emerged as promising alternatives offering significant computational efficiency (guo2016convolutional, ; zhu2019physics, ; geneva2020modeling, ; wandel2021teaching, ; huang2024lordnet, ; zhou2024unisolver, ; alkin2024universal, ; gao2025generative, ). These approaches approximate PDE solution operators, enabling rapid inference once trained. However, they face a fundamental challenge: error accumulation during autoregressive rollouts, where small errors compound exponentially, leading to numerical instability and divergence (Sanchez-Gonzalez2020, ).
Current mitigation strategies, such as injecting random noise during training (Sanchez-Gonzalez2020, ) or multi-step training regimes, often fall short because prediction errors are structured rather than random. Newer approaches like PDE-refiner (lippe_pde-refiner_2023, ) employ generative diffusion models at each step, but introduce substantial computational overhead that may negate the speed advantages of neural solvers.
We propose a fundamentally different approach that directly corrects each prediction using the governing PDE itself. Our key insight is that the PDE residual provides a natural signal for correction. By formulating this as a linear inverse problem based on the Jacobian of the PDE residual, we efficiently project predictions onto the manifold of physically consistent solutions (Figure 1).
Our approach requires no additional training, operates efficiently during inference, and is compatible with any pretrained neural PDE solver. For many PDEs, the Jacobian matrix and its pseudoinverse can be precomputed in an offline warm-up phase, resulting in minimal computational overhead. Even for highly nonlinear PDEs with imperfect Jacobian approximations, the correction significantly improves long-term stability. Our contributions are:
  • A physics-informed correction framework that leverages PDE residuals for stable long-term rollouts without retraining neural models.
  • An efficient caching strategy that precomputes the Jacobian pseudoinverse during an offline phase, reducing computational burden while maintaining accuracy.
  • Demonstration of broad applicability across multiple PDE types (Navier-Stokes, wave equation, Kuramoto-Sivashinsky) and neural architectures (FNO, UNet, ViT), showing consistent improvements in accuracy and stability.
In the following section, we review related work on neural PDE solvers and existing approaches for improving rollout stability, before detailing our physics-informed correction framework in Section 3 and validating its performance across diverse PDE systems in Section 4.
2 Background & Related Work
Neural PDE Solvers.
Neural operators have emerged as powerful surrogates for solving time-dependent PDEs, offering significant computational speedups over traditional numerical methods by learning mappings between function spaces to approximate PDE solution operators.
The Fourier Neural Operator (FNO) (li_fourier_2020, ) established a foundation by performing spectral domain convolutions to capture global dependencies with excellent generalization. The field has since evolved with specialized architectures: message-passing networks (brandstetter_message_2022, ) for complex geometries, Clifford neural networks (brandstetter_clifford_2022, ) for physical invariances, and transformer-based approaches like GNOT (hao_gnot_2023, ), Transolver (wu_transolver_2024, ), and CViT (wang_cvit_2025, ) that combine attention mechanisms with physical priors.
For time-dependent PDEs, these models are typically applied autoregressively, predicting the next state from the current one. While they excel at one-step predictions on in-distribution data, they struggle with long-term stability. Small errors accumulate and amplify over multiple time steps, eventually causing catastrophic divergence – one of the most significant barriers to deploying neural operators in real-world applications.
Strategies for Improving Rollout Stability.
Several approaches have been developed to address this challenge. Sanchez-Gonzalez2020 introduced adversarial training with random noise injection to build robustness against perturbations, while brandstetter_message_2022 proposed multi-step training where loss is computed over prediction sequences, allowing models to compensate for their own errors.
Recent advances leverage generative models: PDE-Refiner (lippe_pde-refiner_2023, ) adapts diffusion models to iteratively denoise predicted states, and hu_wavelet_2024 employs wavelet-based diffusion to improve spectral accuracy. While effective, these methods introduce substantial computational overhead during training or inference – often negating the efficiency advantages of neural surrogates.
Physics-Based Correction Approaches.
An alternative strategy leverages the governing equations to correct predictions without requiring model retraining. cao_residual-based_2023 developed PDE residual minimization methods for Bayesian inference using goal-oriented a-posteriori error estimation (jha_goal-oriented_2022, ), later extended to nonlinear variational problems (jha_residual-based_2024, ). While promising, these approaches often introduce significant computational overhead.
Other research has focused on enforcing specific physical constraints: jiang_enforcing_2020 implemented spectral projection layers for divergence-free conditions in fluid simulations, and duruisseaux_towards_2024 generalized this to broader linear differential constraints. These methods are computationally efficient but limited to specific physical aspects rather than ensuring complete PDE satisfaction.
Our Approach.
Our work bridges these approaches through a correction framework that enforces PDE satisfaction at each timestep via residual minimization. Unlike existing methods, our approach requires no additional training, employs efficient caching strategies to minimize computational overhead, and generalizes across different PDE types and neural architectures. By treating physical consistency as an online correction problem we achieve stable long-term rollouts while preserving the computational advantages of neural PDE solvers.
3 Methodology
Long-term Evolution with Neural PDE Solvers.
Consider time-dependent partial differential equations of the general form:
$S(\mathbf{u},\frac{\partial\mathbf{u}}{\partial t},\frac{\partial^{2}\mathbf{u}}{\partial^{2}t},\frac{\partial\mathbf{u}}{\partial\mathbf{x}},\frac{\partial^{2}\mathbf{u}}{\partial^{2}\mathbf{x}})=0,$ (1)
where $\mathbf{u}$ represents the solution defined on spatial domain $\mathbf{x}\in\mathcal{X}$ and temporal domain $t\in[0,T]$.
Neural PDE solvers approximate the time-evolution operator of such systems. Given a current state $\mathbf{u}(\mathbf{x},t)$, a neural network $\phi_{\theta}$ with parameters $\theta$ predicts the state at the next time step:
$\mathbf{u}(\mathbf{x},t+\Delta t)=\phi_{\theta}(\mathbf{u}(\mathbf{x},t)).$ (2)
For long-term simulations starting from an initial condition $\mathbf{u}_{0}$, the state at time $t$ is obtained through repeated application of $\phi_{\theta}$:
$\mathbf{u}_{t}=\phi_{\theta}(\phi_{\theta}(...\phi_{\theta}(\mathbf{u}_{0})+\epsilon_{0}...)+\epsilon_{t-1}),$ (3)
where $\epsilon_{t}$ represents the prediction error at step $t$.
The Challenge of Error Accumulation.
The fundamental challenge in autoregressive rollouts is that prediction errors compound over time, often leading to numerical instability or complete divergence from the true solution trajectory. Existing approaches attempt to address this problem by either: (1) Enhancing model robustness through specialized training regimes like noise injection or multi-step training, which requires substantial additional training data and computational resources; or (2) Applying post-hoc corrections through computationally expensive denoising procedures that may negate the efficiency advantages of neural solvers.
Both approaches face limitations because prediction errors exhibit complex, structured patterns rather than random noise. These errors depend on the specific state distribution and are difficult to anticipate through data augmentation alone. Moreover, even highly accurate neural operators may eventually diverge during sufficiently long rollouts.
Here we propose an alternative paradigm: a lightweight, physics-informed correction mechanism that operates during inference without requiring model retraining. By explicitly minimizing the PDE residual at each time step, we project predictions back onto the manifold of physically valid solutions, effectively transforming a challenging multi-step prediction problem into a sequence of more manageable one-step predictions. This approach aims to strike an optimal balance between computational efficiency and numerical stability, providing a general solution for accurate long-term simulations across different PDE types and neural architectures.
3.1 The PhysicsCorrect Framework: Linearized PDE Residual Correction
Problem Formulation.
The core idea of our PhysicsCorrect approach is to leverage the governing PDE itself as a form of implicit supervision, correcting neural network predictions to better satisfy the underlying physics. For a state $\mathbf{u}_{t}$ at time $t$, a neural operator produces a prediction $\hat{\mathbf{u}}_{t+1}$ for the next time step. Our goal is to find a correction term $\mathbf{u}_{t+1}^{c}$ such that the corrected prediction $\hat{\mathbf{u}}_{t+1}+\mathbf{u}_{t+1}^{c}$ better approximates the true solution $\mathbf{u}_{t+1}$.
While the ground truth $\mathbf{u}_{t+1}$ is unavailable during inference, we can evaluate how well a candidate solution satisfies the governing equation by computing the PDE residual. For a discretized PDE, this residual $L_{\text{PDE}}(\mathbf{u}_{t},\mathbf{u}_{t+1})$ is obtained by substituting $\mathbf{u}_{t}$ and $\mathbf{u}_{t+1}$ into Equation 1. A physically consistent solution would yield a residual of zero (or for second-order time derivatives, the residual would depend on $\mathbf{u}_{t}$, $\mathbf{u}_{t+1}$, and $\mathbf{u}_{t-1}$).
Efficient Correction via Linear Approximation.
A natural approach would be to directly minimize the PDE residual with respect to the correction term:
$\mathbf{u}_{t+1}^{c*}=\underset{\mathbf{u}_{t+1}^{c}}{\arg\min}\;\|L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1}+\mathbf{u}_{t+1}^{c})\|^{2}.$ (4)
However, directly solving this optimization problem would require iterative methods with adaptive learning rates, introducing significant computational overhead that could negate the efficiency advantages of neural PDE solvers. Instead, we linearize the problem using a first-order Taylor expansion of the residual around the current prediction $\hat{\mathbf{u}}_{t+1}$:
$L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1}+\mathbf{u}_{t+1}^{c})\approx L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})+\frac{\partial L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})}{\partial\hat{\mathbf{u}}_{t+1}}\mathbf{u}_{t+1}^{c}.$ (5)
This approximation is valid when the correction term $\mathbf{u}_{t+1}^{c}$ is sufficiently small, which is generally
Figure 2: Long-term rollout accuracy comparison for the 2D Navier-Stokes benchmark. The baseline neural operator (brown) exhibits error accumulation, while our predictor-corrector approach (blue) maintains stability throughout the simulation, closely matching the performance of idealized one-step rollouts (yellow).
the case when the neural operator has been adequately trained. Setting the linearized residual to zero yields a linear system:
$\frac{\partial L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})}{\partial\hat{\mathbf{u}}_{t+1}}\mathbf{u}_{t+1}^{c}=-L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1}).$ (6)
This can be expressed in the standard form $A\mathbf{x}=\mathbf{b}$, where $A$ is the Jacobian matrix of the PDE residual with respect to $\hat{\mathbf{u}}_{t+1}$, $\mathbf{x}$ is the correction term $\mathbf{u}_{t+1}^{c}$ we seek to determine, and $\mathbf{b}$ is the negative PDE residual $-L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})$. The correction term can then be obtained by solving this linear system, typically using least-squares methods to handle potential ill-conditioning or over-determined systems. Unlike training-based approaches that learn to predict corrections from residuals, our method directly inverts the linearized PDE operator, avoiding the need for additional training data and mitigating potential distribution shifts between training and inference.
Predictor-Corrector Pipeline.
Based on this formulation, we implement PhysicsCorrect as a two-step predictor-corrector pipeline:
  1. 1.
    Prediction Step: The neural operator $\phi_{\theta}$ produces a prediction $\hat{\mathbf{u}}_{t+1}=\phi_{\theta}(\mathbf{u}_{t})$.
  2. 2.
    Correction Step: We solve the linear system in Equation 6 to obtain the correction term $\mathbf{u}_{t+1}^{c}$ and compute the corrected prediction $\tilde{\mathbf{u}}_{t+1}=\hat{\mathbf{u}}_{t+1}+\mathbf{u}_{t+1}^{c}$.
This corrected state $\tilde{\mathbf{u}}_{t+1}$ then serves as input for the next prediction step. By ensuring that each state better satisfies the underlying PDE, we significantly reduce error accumulation during autoregressive rollouts. As shown in Figure 2, this approach maintains stability and accuracy over hundreds of time steps when applied to the 2D Navier-Stokes equation, while the baseline prediction eventually diverges due to accumulated errors. The figure also illustrates how our method’s performance closely approaches that of idealized one-step rollouts, effectively transforming a challenging multi-step prediction problem into a sequence of more manageable one-step predictions.
Our formulation assumes that the neural operator’s prediction provides a good initial approximation, allowing the linearization to be effective. This assumption generally holds for well-trained models but may be challenged in highly chaotic systems or with poor initialization. The effectiveness of our approach across different PDEs and neural architectures is demonstrated empirically in Section 4.
Efficient Implementation via Jacobian Caching.
While the physics-informed correction significantly improves rollout stability, a naive implementation would introduce substantial computational overhead from two expensive operations at each time step: (1) evaluating the Jacobian matrix, and (2) solving the resulting least-squares problem. The Jacobian evaluation requires computing gradients between two spatial fields of dimension $M\times N$, where $M$ and $N$ represent the height and width of the domain. This operation, typically performed through automatic differentiation, scales poorly with increasing resolution. Similarly, solving the least-squares problem via singular value decomposition is numerically stable but computationally intensive.
We introduce a key optimization that dramatically reduces this computational burden. For many time-dependent PDEs, we observe that the Jacobian matrix remains constant across different time steps and initial conditions when the residual $L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})$ is linear with respect to $\hat{\mathbf{u}}_{t+1}$. This property allows us to precompute the Jacobian matrix once during an offline warm-up phase, along with its Moore-Penrose pseudoinverse $A^{\dagger}$. During inference, we can then directly compute the correction as $\mathbf{u}^{c}_{t+1}=A^{\dagger}\mathbf{b}$, where $\mathbf{b}=-L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})$. This approach effectively transforms the correction step from an expensive numerical operation to a simple matrix multiplication, resulting in minimal computational overhead during rollout.
For this caching strategy to be effective, the PDE residual must maintain linearity with respect to $\hat{\mathbf{u}}_{t+1}$. To satisfy this requirement while preserving numerical stability, we employ a semi-implicit discretization scheme that treats linear terms (e.g., diffusion) implicitly to ensure stability, while handling nonlinear terms (e.g., advection) explicitly to preserve Jacobian constancy. For example, in the 2D Navier-Stokes equations, we implement a Crank-Nicolson scheme with implicit diffusion and explicit advection terms (Appendix A). This formulation ensures that the Jacobian remains constant across all prediction steps and initial conditions, while maintaining adequate numerical stability.
As demonstrated in Figure 3, our caching strategy achieves accuracy comparable to the non-cached version while reducing computational cost by approximately 160x. The precomputation phase for a 64x64 resolution grid requires only 8.74 seconds, after which the correction adds minimal overhead to the neural operator’s inference time (0.90 vs. 0.69 seconds for 200 time steps). Importantly, even for chaotic systems where the semi-implicit discretization yields an approximate Jacobian, the correction still substantially improves rollout stability. We observe that this approach works effectively even when applying it to systems with nonlinear residual terms, as we demonstrate with the Kuramoto-Sivashinsky equation in Section 4.3.
Figure 3: PhysicsCorrect’s caching strategy efficiency on 2D Navier-Stokes. Left: Relative $L_{2}$ error vs. reference solutions over 200 time steps. Right: PDE residual magnitude per step. While the baseline (brown) shows increasing error and residual, both uncached (yellow) and Jacobian-cached (blue) corrections maintain low values. Pseudoinverse caching (red) preserves performance while reducing computational cost by 163x.
4 Experiments
We evaluate the PhysicsCorrect framework on three representative PDE systems that vary in complexity, dimensionality, and dynamical behavior: the 2D Navier-Stokes equations (incompressible fluid flow), the 2D wave equation (second-order hyperbolic PDE), and the 1D Kuramoto-Sivashinsky equation (fourth-order nonlinear PDE exhibiting chaotic behavior).
For each system, we test our approach with three neural network architectures: Fourier Neural Operator (FNO) (li_fourier_2020, ), UNet (Wandel2020, ), and Vision Transformer (ViT) (dosovitskiy_image_2020, ). All models are trained to generalize across different initial conditions using the Adam optimizer and L1 loss. Performance is measured by the relative L2 error between predictions and high-fidelity numerical solutions.
Our experiments are designed to address three key questions: whether the correction framework improves the stability and accuracy of long-term rollouts across different neural architectures; how effective the caching strategy is in maintaining accuracy while reducing computational overhead; and how the correction performs on systems with different levels of nonlinearity and chaotic behavior. Network architecture details and training parameters are provided in Appendix B.
4.1 2D Navier-Stokes equation
We first evaluate our approach on the 2D incompressible Navier-Stokes equation with a Reynolds number of 1,000 and forcing term $f(x,y)=0.1\sin(2\pi(x+y))+\cos(2\pi(x+y))$ (Appendix B.1). This system exhibits complex vorticity dynamics and is widely used as a benchmark for fluid simulation.
Experimental Setup.
We train our models on a dataset of 1,000 trajectories generated from Gaussian random initial vorticity fields $\omega_{0}\sim\mathcal{N}\left(0,8^{3}(-\Delta+64I)^{-4.0}\right)$, simulated on a $64\times 64$ grid with a time step of 0.01. For the PDE residual formulation, we employ a semi-implicit Crank-Nicolson scheme with explicit advection and implicit diffusion terms. We evaluate performance on 64 test trajectories, each simulated for 1,000 time steps.
One-Step Correction Performance.
Figure 4 demonstrates the effect of our correction on a single prediction step. The baseline FNO prediction contains small but structured errors that our method effectively eliminates. The correction reduces the relative L2 error by an order of magnitude (from 3.3e-5 to 5.5e-6) while simultaneously reducing the PDE residual to near zero, indicating that the corrected state closely satisfies the governing equation.
Long-Term Rollout Stability.
Figure 5 shows the long-term rollout performance across different neural architectures (FNO, UNet, and ViT). All baseline models exhibit error accumulation that eventually leads to complete divergence from the reference solution. In contrast, models augmented with our physics-informed corrector maintain stable and accurate predictions throughout the entire 1,000-step simulation, regardless of the underlying architecture. The PDE residual (shown in Appendix B.1) remains consistently low for all corrected models, confirming that our approach enforces physical consistency at each time step and prevents error accumulation.
The results demonstrate that our correction framework effectively transforms inherently unstable neural PDE solvers into stable simulation tools without requiring architecture-specific modifications or additional training. This universality is particularly valuable as it allows practitioners to leverage any pre-trained neural operator while ensuring physical consistency and long-term stability.
Figure 4: One-step correction on the 2D Navier-Stokes equation. From left to right: ground truth solution from numerical simulation, prediction error of baseline FNO, prediction error after our correction, PDE residual of baseline prediction, and PDE residual after correction. Note the significant reduction in both error magnitude (10× improvement) and PDE residual (100× improvement), demonstrating that our correction effectively projects predictions onto the manifold of physically consistent solutions.
Figure 5: Performance comparison of our physics-informed correction approach across different PDE systems and neural architectures. Left axis (bars): Relative L2 error of baseline models (lighter colors) versus corrected models (darker colors) for Navier-Stokes (NS), wave equation, and Kuramoto-Sivashinsky (KS) equations at the final state after long rollouts (1000 time step rollout for NS and KS equations; 100 time steps for wave equation). Right axis (line): Relative computational cost of the corrected approach compared to baseline. The correction framework consistently reduces error across all PDEs and architectures with minimal computational overhead. Detailed rollout histories are provided in the Appendix.
Understanding the Limits of Residual-Based Correction.
Even with perfect numerical methods, discretization introduces small but non-zero PDE residuals in reference solutions, as shown in the leftmost panel of Figure 6. Since our method targets zero residual, this creates a fundamental discrepancy – we optimize toward a slightly different objective than the true numerical solution, introducing small inherent errors (fourth panel of Figure 6).
To investigate this limitation, we conducted an idealized experiment where we subtract the reference solution’s residual from the prediction’s residual before correction. This adjustment aligns our optimization target precisely with the numerical reference, resulting in significantly improved accuracy (relative L2 error reduced from 3.3e-5 to 6.1e-7) as shown in the rightmost panel of Figure 6. While this approach is impractical for real applications where reference solutions are unavailable, it demonstrates the theoretical upper bound of our method’s performance.
This experiment yields two important insights: first, the quality of PDE discretization directly impacts correction accuracy; and second, even with standard discretization, our method achieves substantial error reduction (approximately 85%) relative to the theoretical optimum. These findings suggest that using finer discretization schemes for residual computation could further improve correction performance in practice.
Figure 6: The visualization of one test sample: PDE residual using (from left to right) numerical reference result, corrected prediction, and the corrected prediction with reference PDE residual, the errors of corrected prediction, and the errors of corrected prediction with reference PDE residual.
4.2 2D wave equation
We next evaluate our approach on the 2D wave equation – a second-order linear PDE that models various physical phenomena including mechanical waves, electromagnetic waves, and seismic propagation (Appendix B.2).
Experimental Setup.
We generate data on a $128\times 128$ grid using 512 Gaussian random fields as initial conditions with periodic boundaries. Training data consists of the first 10 time steps (recorded at intervals of $\Delta t=10^{-2}$), while testing evaluates generalization over 100 time steps. For residual computation, we employ an implicit scheme with central finite differences for the time derivatives.
Neural Architecture Considerations.
An interesting finding emerged during our wave equation experiments: standard first-order residual prediction (predicting $\mathbf{u}_{t+1}-\mathbf{u}_{t}$) consistently failed to capture the essential physics. To resolve this we revisited the problem through the lens of second-order dynamics, training networks to predict $\delta\mathbf{u}_{t}=\mathbf{u}_{t+1}+\mathbf{u}_{t-1}-2\mathbf{u}_{t}$ instead. This formulation resonates with the oscillatory nature of wave phenomena, providing a significantly more stable formulation. This insight highlights how the representation of physical dynamics can fundamentally shape neural network performance, even before correction mechanisms are applied.
Results.
Figure 5 shows the long-term rollout performance of different neural architectures with and without our correction framework. Even with the improved second-order formulation, baseline models (particularly ViT and UNet) still exhibit error growth over time. Our physics-informed corrector consistently enhances prediction accuracy across all architectures, maintaining low PDE residuals throughout the simulation.
The wave equation results demonstrate our method’s effectiveness on linear PDEs with oscillatory dynamics. Interestingly, the performance improvement is more pronounced for architectures that struggle more with the baseline formulation (ViT and UNet), suggesting that our correction approach can help compensate for architecture-specific weaknesses in capturing certain physical dynamics.
4.3 Kuramoto-Sivashinsky equation
Our final and most challenging test case is the Kuramoto–Sivashinsky (KS) equation – a fourth-order nonlinear PDE that exhibits chaotic dynamics (Appendix B.3). This system tests our method’s effectiveness on strongly nonlinear, chaotic dynamics where prediction errors can amplify rapidly and where linearization approximations are most severely challenged.
Experimental Setup.
We simulate the KS equation on a spatial domain $[0,64]$ with a resolution of 512 points, focusing on the chaotic regime. Starting from $v(x,50)$, we generate 512 trajectories for training (first 500 steps) and 64 trajectories for testing (full 1000 steps) using a spectral method with temporal step size 0.05 (dresdner_learning_2023, ).
Challenges with Chaotic Systems.
The KS equation presents unique challenges for our correction approach. The standard strategy of using semi-implicit discretization (implicit for linear terms, explicit for nonlinear terms) to obtain a constant Jacobian proves inadequate due to the equation’s strong nonlinearity and chaotic behavior. A fully implicit discretization would provide better residual definition but would require re-computing the Jacobian and its pseudoinverse at each time step, negating the computational advantages of our caching strategy.
Effectiveness of Approximate Correction.
Surprisingly, as shown in Figure 5, our approach with cached pseudoinverse still provides significant stability improvements despite using an approximation of the true Jacobian. Figure 7(a) shows a qualitative comparison between baseline FNO predictions (top) and corrected predictions (bottom), demonstrating that our method successfully maintains the complex spatiotemporal patterns of the KS equation over long rollouts.
To further investigate the impact of Jacobian approximation in chaotic systems, we conducted additional experiments varying the frequency of Jacobian updates, as shown in Figure 7(b). The results reveal that recalculating the Jacobian every 3-10 time steps provides minimal improvement over the fully cached approach (updating only once at initialization). This suggests that for the KS equation, the primary benefit comes from the initial projection toward the physically consistent manifold rather than from having a perfectly accurate Jacobian at each step.
These findings lead to an important insight: accurate definition of the PDE residual is more critical than perfect estimation of the Jacobian pseudoinverse. Even with an approximate linear correction, projecting predictions toward the physically consistent manifold provides sufficient regularization to prevent error accumulation. This observation is particularly significant for chaotic systems, where traditional methods often struggle to maintain long-term stability, and offers a favorable trade-off between computational efficiency and prediction accuracy.
Figure 7: Performance of PhysicsCorrect on the chaotic Kuramoto-Sivashinsky equation. (a) Spatiotemporal evolution over 1000 time steps comparing ground truth, baseline FNO prediction, and error magnitude for uncorrected (top) and corrected (bottom) predictions. (b) Impact of Jacobian update frequency on KS equation prediction error over 1000 steps. Periodic recomputation offers minimal improvement over the fully cached approach, while semi-implicit discretization for both Jacobian and residual shows notably higher errors.
5 Discussion
Summary of Contributions.
We introduced PhysicsCorrect, a training-free, physics-informed correction framework that significantly enhances neural PDE solver stability during long-term rollouts. Our approach works with any pretrained neural operator with minimal computational overhead through efficient caching. Across Navier-Stokes, wave equation, and Kuramoto-Sivashinsky systems, our method reduces prediction errors by 1-2 orders of magnitude compared to baselines, with consistent benefits across FNO, UNet, and ViT architectures. By enforcing physical consistency at each step, we transform multi-step rollouts into sequences of well-conditioned one-step predictions. Remarkably, even with approximate linearization and cached Jacobians, the method substantially improves stability in chaotic systems, demonstrating robust practical utility.
Limitations & Future Work.
Despite its effectiveness, PhysicsCorrect faces three key limitations that suggest directions for future work: (1) computational scalability to high-resolution simulations, as Figure 8 demonstrates quadratic scaling of time and memory requirements with resolution, which could be mitigated by inverting the Jacobian approximately using iterative methods that only require Jacobian-vector products; (2) discretization-induced errors creating an inherent gap between our correction target and true numerical references, potentially improvable through higher-order discretization schemes; and (3) potential breakdown of linearization approximations in extreme chaotic systems, which might be addressed by hybrid approaches combining efficient linear correction with occasional nonlinear optimization steps for systems with extreme sensitivity or strong nonlinearities.
Figure 8: PhysicsCorrect’s computational scaling with grid resolution for 2D Navier-Stokes. Plots show computation time (left), GPU memory usage (center), and peak memory consumption (right). All metrics scale quadratically with resolution, highlighting memory and compute requirements as the primary limitation for high-resolution applications.
Conclusions.
This work demonstrates that enforcing physical consistency through direct projection onto the manifold of valid solutions provides a powerful, computation-efficient approach to stabilizing neural PDE solvers without requiring additional training or expensive denoising. The method’s simplicity, generality, and efficiency make it immediately applicable across scientific and engineering applications, effectively bridging the gap between deep learning’s computational advantages and the physical fidelity demanded by practical applications.
Appendix A PDE Residuals Approximation
We detail the numerical discretization employed in this work for defining the PDE residuals, along with relevant implementation considerations. In all experiments, the computational domain $\Omega\subseteq\mathbb{R}^{2}$ is discretized into an $n\times n$ uniform Cartesian grid with mesh spacing $\Delta$. Any function defined on $\Omega$ is then approximated by its values at the grid points. Specifically, we denote the discrete approximation of a function function $u(x,y)$ at the gird point $(x_{i},y_{j})$ by $u_{i,j}$, where $x_{i}=x_{0}+i\Delta$ and $y_{j}=y_{0}+j\Delta$ for $i=0,\cdots,n-1$ and $j=0,\cdots,n-1$. For time-dependent variables, we discretize the time interval $[0,T]$ with a uniform time step $\Delta t$, and denote a discrete function $u$ at the spatial gird point $(i,j)$ and $t$-th time step by $u_{i,j}^{t}$.
2D Navier-Stokes equation. Considering incompressible two-dimensional Navier-Stokes equations with periodic boundary conditions as follows:
$\begin{gathered}\frac{\partial\omega}{\partial t}=-\frac{\partial\psi}{\partial y}\frac{\partial\omega}{\partial x}+\frac{\partial\psi}{\partial x}\frac{\partial\omega}{\partial y}+\frac{1}{\operatorname{Re}}\left(\frac{\partial^{2}\omega}{\partial x^{2}}+\frac{\partial^{2}\omega}{\partial y^{2}}\right)+f(x,y),\\ \frac{\partial^{2}\psi}{\partial x^{2}}+\frac{\partial^{2}\psi}{\partial y^{2}}=-\omega,\end{gathered}$ (7)
in which both vorticity function $\omega$ and steam function $\psi$ are time-dependent, and $f$ is the forcing term. Given the stream function $\psi^{t}_{i,j}$ at time step $t$ and a one-step prediction $\hat{\psi}^{t+1}_{i,j}$ for $i=1,\cdots,n$ and $j=1,\cdots,n$, we first apply central difference based differential operator to compute the corresponding vorticities $\omega^{t}_{i,j}$ and $\hat{\omega}^{t+1}_{i,j}$. To integrate the governing equation (Equation 7), we adopt an implicit scheme for the diffusion term to ensure numerical stability, while an explicit scheme for the advection part to avoid non-linearity. Using a forward Euler method in time, the PDE residual for Equation 7 used to approximate the Jacobian matrix is then defined as follows,
$L_{\text{PDE}}=\sum_{i,j}^{n\times n}\Big{|}\frac{1}{\Delta t}(-\hat{\omega}^{t+1}_{i,j}+\omega^{t}_{i,j})-\frac{1}{4\Delta^{2}}\left[\left(\psi_{i,j+1}^{t}-\psi_{i,j-1}^{t}\right)\right.\left.\left(\omega_{i+1,j}^{t}-\omega_{i-1,j}^{t}\right)\right.$ (8)
$\left.-\left(\psi_{i+1,j}^{t}-\psi_{i-1,j}^{t}\right)\left(\omega_{i,j+1}^{t}-\omega_{i,j-1}^{t}\right)\right]$
$+0.5\frac{1}{\text{Re}\Delta^{2}}\left(\omega_{i+1,j}^{t}+\omega_{i-1,j}^{t}+\omega_{i,j+1}^{t}+\omega_{i,j-1}^{t}-4\omega_{i,j}^{t}\right.$
$\left.+\hat{\omega}_{i+1,j}^{t+1}+\hat{\omega}_{i-1,j}^{t+1}+\hat{\omega}_{i,j+1}^{t+1}+\hat{\omega}_{i,j-1}^{t+1}-4\hat{\omega}_{i,j}^{t+1}\right)+f_{i,j}\Big{|}$
We can see that $\frac{\partial L_{\text{PDE}}}{\partial\hat{\psi}^{t+1}}$ depends only on the current state $\psi^{t}$.
2D Wave Equation. The two-dimensional wave equation with constant wave speed is formulated as
$\frac{\partial^{2}u}{\partial t^{2}}=c^{2}(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}),$ (9)
where the displacement $u$ is time-dependent and $c$ is the constant wave speed. We use the second-order central difference to approximate the Laplace operator here, so we have
$\frac{\partial^{2}u}{\partial t^{2}}\Big{|}_{i,j}=c^{2}\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right)_{i,j}=$ $\frac{u_{i-1,j}-2u_{i,j}+u_{i+1,j}}{\Delta^{2}}+\frac{u_{i,j-1}-2u_{i,j}+u_{i,j+1}}{\Delta^{2}},$ (10)
at each grid point $(x_{i},y_{j})$ for $i=1,\cdots,n$ and $j=1,\cdots,n$. Since we did not have fine sampling in the temporal direction, we use an implicit scheme and central finite difference for the second-order time derivative. Given the displacement function $u^{t-1}_{i,j}$ at time $t-1$ and $u^{t}_{i,j}$ at time $t$, and the one-step prediction $\hat{u}^{t+1}_{i,j}$ at time $t+1$, the discretized PDE residual is represented as:
$L_{\text{PDE}}=\sum_{i,j}^{n\times n}\Big{|}$ $\frac{\hat{u}^{t+1}_{i,j}+u^{t-1}_{i,j}-2u^{t}_{i,j}}{\Delta t^{2}}-$
$\frac{c^{2}}{3}\left(\frac{\partial^{2}u^{t-1}}{\partial x^{2}}+\frac{\partial^{2}u^{t-1}}{\partial y^{2}}+\frac{\partial^{2}u^{t}}{\partial x^{2}}+\frac{\partial^{2}u^{t}}{\partial y^{2}}+\frac{\partial^{2}\hat{u}^{t+1}}{\partial x^{2}}+\frac{\partial^{2}\hat{u}^{t+1}}{\partial y^{2}}\right)_{i,j}\Big{|}.$ (11)
Hence, $\frac{\partial L_{\text{PDE}}}{\partial\hat{u}^{t+1}}$ does not depend on the prediciton $\hat{u}^{t+1}$ itself. Notably, for the wave equation, being a linear system, the PDE residual in Equation 5 does not rely on the assumption of small prediction error, as no Taylor expansion is needed. This further implies that PhysicsCorrect remains robust to prediction errors in linear systems, a phenomenon that will be demonstrated empirically in later sections.
Kuramoto-Sivashinsky Equation. The Kuramoto-Sivashinsky equation with a constant viscosity parameter of 1.0 is as follows:
$v_{t}+v_{xx}+v_{xxxx}+vv_{x}=0.$ (12)
We apply the spectral method to obtain the spatial derivative for different orders. Given $v^{t}$ at $t$-th time step and a one-step prediction $\hat{v}^{t+1}$ at $t+1$-th time step, the PDE residual is defined as follows:
$L_{\text{PDE}}=\sum_{i}^{n}\Big{|}\frac{\hat{v}^{t+1}_{i}-v^{t}_{i}}{\Delta t}+0.5(\frac{\partial^{2}v^{t}}{\partial x^{2}}+\frac{\partial^{4}v^{t}}{\partial x^{4}}+v^{t}\frac{\partial v^{t}}{\partial x}+\frac{\partial^{2}\hat{v}^{t+1}}{\partial x^{2}}+\frac{\partial^{4}\hat{v}^{t+1}}{\partial x^{4}}+\hat{v}^{t+1}\frac{\partial\hat{v}^{t+1}}{\partial x})_{i}\Big{|},$ (13)
where $i$ denotes the spatial points. We simplify this by pre-computing the pseudoinverse of the Jacobian matrix using a semi-implicit discretization for the nonlinear term and an implicit scheme for the linear term, which approximates the time-dependent Jacobian $\frac{\partial L_{\text{PDE}}}{\partial\hat{v}^{t+1}}$. In the subsequent correction steps, all terms in the PDE residual are treated implicitly to further improve the accuracy of the residual formulation.
Appendix B Experimental Details
In this section, we introduce the details of the configurations of different benchmarks and neural network configurations, as well as additional visualization results. The code and data will be available at https://github.com/summerwine668/PhysicsCorrect.
B.1 Navier-Stokes Equation
We generate 1,000 simulation trajectories as training data using a Reynolds number of 1,000 and a forcing term defined as $f(x,y)=0.1\sin(2\pi(x+y))+\cos(2\pi(x+y))$. Each trajectory begins from a Gaussian random vorticity field sampled from the distribution $\mathcal{N}\left(0,8^{3}(-\Delta+64I)^{-4.0}\right)$. An additional 64 trajectories, generated using the same configuration, are used for testing. For training, we include the first 100 time steps of each trajectory, while evaluation measures generalization performance over 1,000 time steps. The neural network is trained to predict the one-step residual $\delta\psi^{t}=\psi^{t+1}-\psi^{t}$. We use an initial learning rate of 1e-3 and a weight decay of 1e-5 for the Adam optimizer, and decay the learning rate by a factor of 0.9 every 5,000 iterations.
Regarding the neural network architectures, we implement the standard Fourier Neural Operator (FNO) following [29], which consists of four Fourier layers. Each layer uses 12 Fourier modes in both the $x$ and $z$ directions, and the width of the feature map is set to 64. The activation function used is the Gaussian Error Linear Unit (GELU) [30]. After the Fourier layers, a linear projection is applied to map the high-dimensional latent representation back to the spatial domain. For the UNet architecture, we adopt the configuration described in [31], with a hidden size of 128. For the Vision Transformer (ViT), we use a patch size of 8 for spatial embedding and stack 4 standard pre-norm transformer blocks. Each block contains four attention heads, with each head having a dimensionality of 128. GELU is also used as the activation function for ViT.
After training for 30,000 iterations with a batch size of 128, we evaluate different baseline models and their corresponding corrected simulation trajectories. Figure 9 presents the relative L2 norm error (computed by dividing the error norm by the ground-truth norm) and the PDE residual histories over the rollout horizon. While all baseline models suffer from error accumulation over time, the proposed corrector yields stable and accurate long-term predictions. The residual histories further confirm that our method consistently enforces physical constraints at each time step.
Figure 9: Long-term rollout performance on the 2D Navier-Stokes equation across different neural architectures. Left: Relative L2 error over 1,000 time steps, showing that baseline models (solid lines) suffer from error accumulation and eventual divergence, while our corrected models (dashed lines) maintain stable and accurate predictions throughout the simulation. Right: PDE residual magnitude, where lower values indicate better satisfaction of the governing equation. Results are averaged over 5 random seeds, with shaded regions showing standard deviation.
B.2 2D Wave Equation
We fixed the wave speed $c$ as 1.0 in our test. We generate Gaussian random fields as initial conditions and then use fourth-order central differences for spatial derivatives and a fourth-order Runge–Kutta (RK4) scheme for time integration, to accurately evolve the displacement field, following [32]. The data are generated on a 128$\times$128 grid with a time step of 1e-4 and recorded every 100 time steps. We sampled 512 trajectories with the first 10 steps (with an interval of 1e-2) for training and another 64 trajectories with all 100 steps for testing. We use an initial learning rate of 1e-3 with a decay factor of 0.6 every 100 epochs.
The FNO configuration for the wave equation is identical to that used for the Navier–Stokes equation. However, the UNet used in this case has a reduced hidden size of 20 and uses LeakyReLU activation function with a negative slope of 0.1. For the ViT, we use a patch size of 4 and increase the embedding dimension to 256. The model consists of 3 transformer blocks, each with eight attention heads and a head dimension of 256.
As mentioned earlier, the choice between predicting the first-order residual (predicting $u^{t+1}-u^{t}$) or the second-order residual (predicting $\delta u^{t}=u^{t+1}+u^{t-1}-2u^{t}$) results in noticeably different long-rollout behaviors, significantly affecting rollout stability and error accumulation. Here, we visualize the result of a representative test sample using FNO, as shown in Figure 10. The baseline rollout with first-order residual prediction diverges rapidly, whereas the rollout with the proposed corrector remains stable. When the network is trained to predict the second-order residual, both the baseline model and the corrected version exhibit stable performance. However, the improvements in this case are marginal for two reasons: (1) the baseline model already performs well and closely satisfies the PDE residual; and (2) discretization-induced errors, as discussed in Figure 6, impose an inherent limitation on correction accuracy.
Figure 10: Comparison of rollout predictions for a representative test sample using different network output strategies and correction methods. The top row shows the ground-truth evolution from $t=0$ to $t=100$. The second row shows the result of first-order residual prediction (trained with full horizon), which rapidly diverges. The third row applies our proposed correction to the first-order residual prediction, significantly improving stability. The fourth and fifth rows show the predictions using second-order residuals (trained with the first 10 time steps), with and without correction, both yielding accurate and stable long-term rollouts.
Next, we train different baseline models to predict the second-order residual for 1,000 epochs with a batch size of 128. Figure 11 shows the relative L2 error and PDE residual histories over 100 time steps. While the improvement of our approach on the FNO baseline is modest, significant gains are observed for the UNet and ViT baselines. Once again, the PDE residual plots indicate that our approach consistently enforces physical fidelity at each time step.
Figure 11: Long-term rollout performance on the 2D wave equation using second-order residual prediction. Left: Relative L2 error over 100 time steps for different neural architectures with (dashed lines) and without (solid lines) our physics-informed corrector. Right: PDE residual magnitude during rollout. While second-order formulation already provides reasonable baseline performance (especially for FNO), our correction framework consistently reduces both prediction error and PDE residual across all architectures. Results are averaged over 5 random seeds with shaded regions showing standard deviation.
B.3 Kuranmoto-Sivashinksy Equation
We aim to solve the Equation 12 on a temporal domain $[0,100]$ and a spatial domain $[0,64]$ while focusing on the chaotic part. The equation is numerically solved using the spectral method [28], with a spatial interval of 0.125 and a temporal interval of 0.05, for data generation. We begin from $v(x,50)$ and train a one-step predictor using neural networks to learn the residual $\delta v^{t}=v^{t+1}-v^{t}$. The training set includes 512 trajectories, using only the first 500 time steps of each, while the test set contains 64 additional trajectories, each with the full 1,000-step rollout. All models are trained with a learning rate of 1e-3, a batch size of 256, and a total of 3,000 training epochs.
For the network configurations, we use a one-dimensional variant of the FNO (FNO-1D), adapted from its 2D counterpart by replacing all 2D convolutional operations with 1D convolutions. The number of Fourier modes is set to 16, and the feature map width is 64. GELU is used as the activation function. The 1D UNet is similarly adapted from the 2D UNet by substituting all 2D convolutional layers with 1D versions. The hidden size is set to 64. We use the LeakyReLU activation function with a negative slope of 0.1, along with Group Normalization [33]. For the 1D ViT, we use a patch size of 4 and an embedding dimension of 768. The model consists of 4 transformer blocks, each containing 16 attention heads with a head dimension of 768. GELU is employed as the activation function throughout the model.
For the corrector, we employ a semi-implicit scheme to precompute the pseudoinverse of the Jacobian matrix, and an implicit scheme for evaluating the PDE residual. Figure 12 presents the relative L2 errors of the baseline models and their corrected counterparts, along with the corresponding PDE residual histories. We observe that our approach consistently improves the rollout performance across all baseline models while maintaining physical consistency at each time step. While applying an implicit scheme for both the pseudoinverse computation and PDE residual evaluation can further improve rollout accuracy (Figure 7b), it incurs substantial computational overhead. Our approach strikes a balance between accuracy and efficiency.
Figure 12: Long-term rollout performance on the KS equation across different neural architectures. Left: Relative L2 error over 1,000 time steps, showing that baseline models (solid lines) suffer from error accumulation and eventual divergence, while our corrected models (dashed lines) maintain relatively stable and accurate predictions throughout the simulation. Right: PDE residual magnitude, where lower values indicate better satisfaction of the governing equation. Results are averaged over 5 random seeds, with shaded regions showing standard deviation.
B.4 Additional Implementation details
Compute Resources. All training is performed on a single NVIDIA A6000 GPU. For the Navier-Stokes equation, training for 30,000 steps with a batch size of 64 takes roughly 60, 12, and 170 minutes using FNO, U-Net, and ViT, respectively. For the wave equation, training for 1,000 epochs takes roughly 29, 6, and 65 minutes using FNO, U-Net, and ViT, respectively. For the KS equation, training for 3,000 epochs takes roughly 100 minutes using FNO, 70 minutes using U-Net, and approximately 5 hours using ViT.
Negligible Additional Inference Time. We also evaluate the inference time of both the baseline models and their counterparts augmented with our proposed correctors. As shown in Figure 5, the ratio between the inference time of the baseline with corrector and that of the baseline model is close to 1.0. Although the Navier–Stokes case with the ViT baseline shows a slightly higher ratio, due to the lightweight nature of the baseline, the additional inference time introduced by our corrector remains negligible, particularly for larger-scale problems and more complex neural networks.
B.5 Sensitivity Analysis: Error Structure and Magnitude Limits
To characterize the operational bounds of PhysicsCorrect, we conducted a systematic sensitivity analysis by artificially corrupting ground truth solutions with controlled error patterns and magnitudes. This approach enables evaluation of corrector robustness under varying error conditions that may arise in practice.
Experimental Protocol.
We simulate different types of prediction errors by adding two distinct noise patterns to reference solutions: uncorrelated Gaussian noise representing random, spatially independent perturbations, and correlated Gaussian random fields representing structured, spatially coherent perturbations, with controlled magnitude (relative L2 error ranging from $10^{-3}$ to 1.0). These corrupted states serve as surrogate neural network predictions, enabling systematic evaluation of corrector performance across different error characteristics.
Linear versus Nonlinear PDE Response.
Figure 13 illustrates the post-correction relative L2 error achieved by the proposed PhysicsCorrect method versus the pre-correction error across different PDE systems.
Figure 13: Post-correction relative L2 error versus pre-correction (one-step) relative L2 error across three PDE systems: Navier–Stokes (NS), Wave, and Kuramoto–Sivashinsky (KS). The x-axis represents the artificially injected corruption level before correction, while the y-axis shows the resulting error after applying the PhysicsCorrect method. Results are reported for both unstructured Gaussian noise and structured Gaussian random field perturbations. The dashed line indicates the uncorrected pre-correction error (i.e., the identity line), serving as a reference. Correction consistently reduces error magnitude, with stronger suppression observed for structured noise.
The diagonal dashed line serves as a reference indicating no correction. Across all systems, our method significantly reduces the prediction error compared to the uncorrected baseline, particularly when the corruption is structured (Gaussian random field). The analysis reveals a fundamental distinction between linear and nonlinear systems. For the wave equation, being a linear PDE with exact Jacobian computation, the corrector maintains effectiveness across the entire tested error range, indicating that the PhysicsCorrect module has likely reached the discretization error floor. The linear nature of the governing equation ensures that our correction mechanism remains theoretically sound regardless of error magnitude because the PDE residual is linear to the prediction, where Equation 5 does not rely on the assumption of small relative error on the prediction for Taylor expansion.
In contrast, the nonlinear Navier-Stokes and Kuramoto-Sivashinsky equations exhibit threshold behavior. Within moderate error levels, correction accuracy remains high, demonstrating the robustness of PhysicsCorrect. However, beyond certain thresholds, where the assumption of local approximation underlying the Taylor expansion no longer holds, the corrected relative error begins to increase with the pre-correction error. This degradation is further attributed to the semi-implicit discretization schemes used for nonlinear PDEs, which yield approximate rather than exact Jacobians. As a result, the linearization becomes less valid under large perturbations. Nevertheless, the results demonstrate that even when the corrected predictions are not back onto the manifold of physically consistent solutions, PhysicsCorrect still offers meaningful accuracy improvements.
Error Structure Dependencies.
Correlated errors consistently prove more challenging for the corrector than uncorrelated noise, particularly in nonlinear chaotic systems (Rightmost panel of Figure 13). This observation is significant because neural network prediction errors typically exhibit spatial structure rather than random patterns, suggesting that real-world performance may be more constrained than random noise analysis would indicate.
Practical Implications.
These findings establish that PhysicsCorrect’s effectiveness is bounded by the validity of the underlying linearization approximation. For linear PDEs, this bound is theoretical, while for nonlinear systems, practical thresholds exist beyond which correction degrades. Users should therefore validate their neural operator’s typical error characteristics against these bounds before deployment, and consider implementing error monitoring to detect when correction should be disabled since the original prediction fails.